Find the value of x: log 361/169 =2
x
± 13/19
20/13
19/13
±19/13
±14/19
13/19
Please help to solve.

Answer by nerdybill(7003)

logx(361/169) =2
361/169 = x^2
taking the square root of both sides:
(+-) 19/13 = x